There are 30 students in the group. The marks for the control test were the following: 8 students received “excellent” marks, 12 students - “good” marks, 8 students – “satisfactory”. What is the probability that 3 students invited for answer have “good” marks?


1
Expert's answer
2021-09-09T07:54:26-0400

Given:

n=30

number of students with good marks = 12


when r items are chosen from n items , use combination formula :

"nc_{r} == \frac{n!}{r!(n-r)!}"


P(three students invites for answer have "good marks")= "\frac{12C_{3}}{30C_{3}}" = "\frac{220}{4060}= 0.0542"